### v3

@@ 5 Non-local regime
In this section, we present the result covering the non-local regime, i.e., the regime when one of the $k$ color classes is far from being quasirandom with density $1/k$. We start with the lemma that informally says that if the density of a graph $H$ in a given graphon is small, then the graphon contains a large sparse part. Since it is folklore that complete graphs satisfy the Kohayakawa-Nagle-Rödl-Schacht Conjecture (Conjecture 25 in Section 7), a simple proof of this fact can be found in [31, Theorem 4.2], the next lemma holds for $H$ being a complete graph and so it holds for all graphs $H$. We include a self-contained proof in Appendix for completeness.
@@ 7 Conclusion
Sidorenko’s Conjecture would imply that Conjecture 25 holds for all bipartite graphs. Among non-bipartite graphs, Conjecture 25 is known to hold only for some specific families [29, 36, 31], e.g, complete multipartite graphs, unicyclic graphs, cycles with a single chord, and graphs obtained by gluing complete multipartite graphs in a tree-like way. Theorems 19 and 22 imply that for every graph $G$ with girth at least $50$, there exists a connected rooted graph $H^{\bullet}$ containing $G$ as an induced subgraph and $n_{0}$ such that the graph $H^{\bullet}\oplus K_{m|\ell,n}^{\bullet}$ satisfies Conjecture 25 for all even integers $m,n,\ell\geq n_{0}$ such that $m$ is divisible by $5$ and $n+e\left(H\right)\leq\ell\leq mn/4$.

### v5

@@ 5 Non-local regime
In this section, we present the result covering the non-local regime, i.e., the regime when one of the $k$ color classes is far from being quasirandom with density $1/k$. We start with the lemma that informally says that if the density of a graph $H$ in a given graphon is small, then the graphon contains a large sparse part. Since it is folklore that complete graphs satisfy the Kohayakawa-Nagle-Rödl-Schacht Conjecture (Conjecture 25 in Section 7), a simple proof of this fact can be found in [32, Theorem 4.2], the next lemma holds for $H$ being a complete graph and so it holds for all graphs $H$. We include a self-contained proof in Appendix for completeness.
@@ 7 Conclusion
Sidorenko’s Conjecture would imply that Conjecture 25 holds for all bipartite graphs. Among non-bipartite graphs, Conjecture 25 is known to hold only for some specific families [30, 37, 32, 3], e.g, complete multipartite graphs, unicyclic graphs, cycles with a single chord, graphs obtained by gluing complete multipartite graphs in a tree-like way, and graphs obtained by gluing along independent sets in highly symmetric ways. Theorems 19 and 22 imply that for every graph $G$ with girth at least $50$, there exists a connected rooted graph $H^{\bullet}$ containing $G$ as an induced subgraph and $n_{0}$ such that the graph $H^{\bullet}\oplus K_{m|\ell,n}^{\bullet}$ satisfies Conjecture 25 for all even integers $m,n,\ell\geq n_{0}$ such that $m$ is divisible by $5$ and $n+e\left(H\right)\leq\ell\leq mn/4$.